OpenStudy (anonymous):

When a train is travelling along a straight track at 2 m/s, it begins to accelerate at a = (60v^(-4)) m/s^2, where v is in m/s. Determine its velocity v and its position 3 s after the acceleration.

OpenStudy (lastdaywork):

a = 60v^(-4) dv/dt = 60v^(-4) \[\int\limits v^4 dv = 60 \int\limits dt + C\] You can calculate C by using v = 2 at t = 0 Similarly find distance as a function of time. Finally put t = 3 to get the corresponding velocity and position.

OpenStudy (anonymous):

Thanks! It's actually quite easy!

OpenStudy (lastdaywork):

You're welcome :)